OpenStudy (alex6799):

Simplify the product using the distributive property. (-4h+2)(3h+7)

OpenStudy (anonymous):

-12h^2-22h+14 What you need to do is factor. -4h*3h=-12h^2 -4h*7=-28h 2*3h=6h 7*2=14 then add -28h+6h=-22h

OpenStudy (alex6799):

so add -28h to 6h and then subtract it from -22h?

OpenStudy (alex6799):

Simplify the product using the distributive property. (-4h +2)(3h +7) A. -12h2 - 22h + 14 B. -12h2 - 34h - 14 C. -12h2 + 34h - 14 D. -12h2 + 22h + 14 the full question with the answer choices ^